# A new economic and financial theory of money

### v6

@@ XIX Electronic currency as a religion
The foundation of religious philosophies have been: (1) the existence of a virtual individual of the collective, that is God, (2) the Golden Rule of the collective, and (3) the antithetical nature of resistivity, such as usury and oppression. These are the foundation of Judaism, Islam, Buddhism, Hinduism, Confucius philosophy, and native African Ubuntu philosophy. It is especially true of native American and Aboriginal philosophy that have a focus on sustainability, that is the common good and future.
@@
Figure 29: Homage being given to Mr. Greed with his pile of money.
It is interesting to note that Lorenzo de’ Medici (Parks, 2005), known as Lorenzo il Magnifico, truly wanted to be a benevolent leader who valued and developed the simple beauty of society, art and science. This was the Renaissance that he inspired. It was unfortunate that he had to contend with the resistance, that is the evil, of the Pazzi and Savonarola. The ultimate irony is that his son Giovanni became Leo X, and his nephew Giulio became Clement VII – the Popes that established the Chest of Indulgences and inspired the Protestant Reformation of Martin Luther. The Medici operated the largest bank for most of the 15th century out of Florence, based on debt financing, but making most of its income from currency exchange futures associated with international trade. It fell victim to the fractional reserve banking risk, caught stealing from the public dowry fund for needed reserves, and went out of business in 1494, shortly after the death of Lorenzo in 1492, being founded in 1397 by Lorenzo’s great grandfather, Giovanni di Bicci.
@@ XXI Discussion and Conclusions
This is a new unified economic theory (Glinsky, 2024d). We mean by “new unified economic theory” is that there is a unification of the monetary, that is how the money supply is determined and controlled, with the fiscal investment policy, that is how investments are made in the economy. Traditionally, the monetary policy has been governed by Monetarism that has been uncoordinated with the fiscal policy. With Monetarism, you have a central bank that is controlling the money supply via a discount or prime rate at which the central bank loans out money. This is Autocratic Capitalism with a central profit or resistivity or interest based monetary policy and control. Traditionally things have been handled from the fiscal investment side with the concepts of Keynesianism. That is where you borrow and spend in order to be able to make the investment by the central government. Since you are borrowing, there is an interest rate associated with the borrowing by the issuing of something like T-bills by the US Government or GILTs by the British Government. This is an autocratic. There is a central capitalistic government. This means the government is profit based with this interest based fiscal policy and control. From the firm’s perspective, the optimization is made in terms of profit. But, these two things, the monetary and fiscal policy, end up being uncoordinated.
@@
This constraint is that all of the returns multiplied by both the network and the savings multipliers ($m_{e}S_{0}/T_{I}\sim 10$) less some of the short-term costs (the ones not covered by short-term returns) be greater than zero. The constraint $\Delta M>0$ is that the increase in the demand for the electronic currency be greater than zero. Just so that it is greater than zero, there will be an increase in the value of the currency and no inflation. The difference between the two constraints is stark – short-term returns less all short-term costs versus all of the returns times $m_{e}S_{0}/T_{I}$ less some of the short-term costs being greater than zero. Much more investment is justified by the $\Delta M>0$ inequality than by the $\text{NPV}>0$ inequality. Looking at the three multipliers in the second line of the expression for NPV in Eq. 157 is very instructive. The term for the DCF, or the upper limit on investment, is the total return $R_{mSR}$ (the upper limit on investment in our theory) times three factors. The first factor is a small factor given by the inverse of the network and savings multipliers $T_{I}/m_{e}S_{0}\sim 0.1$. Remember that these two multipliers come from the fact that an increase of the yearly revenue of a member of the sub-economy is multiplied by both the network effect and the savings effect. The second factor is a term that favors projects with a large profit margin. The third factor is a function that favors projects with a short life time $T_{0}\lesssim 1/\nu$, and quick payback $T_{I}\lesssim 1/\nu$. These are all false incentives, with the first term a false reduction in the potential investment because the network and savings multipliers are not recognized by NPV.
@@
The equity of the business is risk tranched into two securities – the electronic currency (that is, the transactional equity) and the stock (that is, the common equity) of the local economy. The electronic currency is the low risk (that is, small fluctuations), low return, easy to transact (that is, trade) security, and the stock is the high risk (that is large fluctuations), high return, more difficult to transact security. The fluctuations of the electronic currency are further reduced, and the equilibrium stabilized by arbitrage trading of the electronic currency. This arbitrage trading transfers the fluctuations (that is, risk) from the electronic currency to the stock. The fluctuations in the electronic currency are driven by a lack of liquidity in the electronic currency, that is the counter party to the trade not being there at the time of the trade. The arbiter steps in and makes the trade, then makes the opposite trade when the counter party arrives. As the local economy gets larger, the liquidity increases and the required liquid reserve ratio decreases, as one over the square root of the size of the local economy. The arbiter buys for a small amount less than the true currency price and sells for a small amount more than the true currency price. The difficulty is knowing the true currency price. When the arbitrage trading is done according to Eq. 100, the arbiter will always buy low and sell high, creating the financial heat pump described in Sec. VII. The risk tranching is a consequence of the backing (that is, securitization) of the electronic currency by tangible assets whose “replacement” costs can be accounted by transactions (or equivalently the transactional monetary demand given in Eq. 16), and the backing of the stock by predominately intangible assets whose value can not be accounted for (that is, things like know-how, show-how, value-in-use, goodwill, and brand value). It is essential that when these two securities are traded by the public, that the public have complete real-time on-line visibility of the finances of the local economy. This allows the public to correctly value the securities.
@@
The bottom line is that electronic currency leveraged financing allows more effective operation of sub-economies and more appropriate levels of fiscal investment. This is supercharged by approaching monetary and fiscal policy as a question of financial systems control based on innovative genAI methods of system characterization (i.e., risk estimation) and system control. The economy is operated at the point of maximum sustainable virtuous economic activity, that is at the point of economic and social prosperity.
@@
The leader has a choice of two ethos: the aesthetic or the unaesthetic. The choice is between a point of global aesthetic maximum where everyone wins or a point of global aesthetic minimum (the unaesthetic) where only a few people win and everyone else looses. This is a choice of the future or the religious eskaton. There are two ways that a strategic player can play a game against a reactionary opponent. The strategic player can lead his opponent to play the game very poorly leading to quick and decisive defeat of the opponent where the strategic player always wins. This is a choice of the unaesthetic, an ugly boring game with the misery of the opponent maximized. The strategic player will feel that the game is boring and unrewarding, too. In contrast, the strategic player can lead his opponent to playing well resulting in a closely matched interesting game where there is a significant chance that the opponent will win. This is a choice of the aesthetic, a beautiful engaging game with the enjoyment of both participants, as well as the spectators, maximized. This is the beautiful game of football of which Brazilian fans speak. It is not whether one wins or looses, it is how one plays the game. The best football match ever played was the 2022 World Cup final between Argentina and France. It does not matter that Argentina won the match. France could have easily won with a slightly different trajectory of the ball off of the foot of Mbappé late in the match. What made it the most beautiful match ever was both Messi and Mbappé playing the best football of their career against each other. They both elevated the play of each other to a perfectly matched level of performance never seen before.
@@
In summary, this paper presented a new unified economic theory that relaxed the capitalistic diffusive approximation made for control so that there is coordinated, yet distributed, monetary and fiscal (print/grant investment) policy made to maximize GDP, not DCF. This is a socialist, not a capitalistic metric, where the economic system is treated as a web (multiscale graph) of economic communes. The local microeconomics is treated like the global macroeconomics. The viscous forecasting and control is replaced by a Generative Artificial Intelligence (genAI) forecasting and control. The control is distributed (that is, democratic), not centralized (that is autocratic). Monetarism argues for only an autocratic (centralized) capitalistic (profit based) monetary control, and Keynesianism argues for only an autocratic capitalistic fiscal (borrow/spend investment) control. This new theory argues for coordinated democratic socialist monetary and fiscal control. It does have philosophical connections to African Ubuntu, the Economics of Mutuality, Modern Monetary Theory, Marxian Economics and religious philosophy. It also makes the connection between genAI and the physics of economic collective systems. The result is an economic system that maximizes the good of society, not an economic system that maximally exploits society.

### v7

@@ XIX Electronic currency as a religion
The foundation of religious philosophies have been: (1) the existence of a virtual individual of the collective, that is God, (2) the Golden Rule of the collective, and (3) the antithetical nature of resistivity, such as usury (that is interest or riba) and oppression. These are the foundation of Judaism, Islam, Buddhism, Hinduism, Confucius philosophy, and native African Ubuntu philosophy. It is especially true of native American and Aboriginal philosophy that have a focus on sustainability, that is the common good and future.
@@
Figure 29: Homage being given to Mr. Greed with his pile of money. Image generated by OpenAI’s DALL-E.
It is interesting to note that Lorenzo de’ Medici (Parks, 2005), known as Lorenzo il Magnifico (1449-1492), truly wanted to be a benevolent leader who valued and developed the simple beauty of society, art and science. This was the Renaissance that he inspired. It was unfortunate that he had to contend with the resistance, that is the evil, of the Pazzi and Savonarola. The ultimate irony is that his son Giovanni became Leo X, and his nephew Giulio became Clement VII – the Popes that established the Chest of Indulgences and inspired the Protestant Reformation of Martin Luther. The Medici operated the largest bank for most of the 15th century out of Florence, based on debt financing, but making most of its income from currency exchange futures associated with international trade. It fell victim to the fractional reserve banking risk, caught stealing from the public dowry fund for needed reserves, and went out of business in 1494, shortly after the death of Lorenzo, being founded in 1397 by Lorenzo’s great grandfather, Giovanni di Bicci.
@@ XXI Discussion and Conclusions
This is a new unified economic theory (Glinsky, 2024g). We mean by “new unified economic theory” is that there is a unification of the monetary, that is how the money supply is determined and controlled, with the fiscal investment policy, that is how investments are made in the economy. Traditionally, the monetary policy has been governed by Monetarism (Fisher, 1911; Friedman, 1959, 1962; Desai, 1981) that has been uncoordinated with the fiscal policy. With Monetarism, you have a central bank that is controlling the money supply via a discount or prime rate at which the central bank loans out money. This is Autocratic Capitalism with a central profit or resistivity or interest based monetary policy and control. Traditionally things have been handled from the fiscal investment side with the concepts of Keynesianism (Keynes, 1936). That is where you borrow and spend in order to be able to make the investment by the central government. Since you are borrowing, there is an interest rate associated with the borrowing by the issuing of something like T-bills by the US Government or GILTs by the British Government. This is an autocratic. There is a central capitalistic government. This means the government is profit based with this interest based fiscal policy and control. From the firm’s perspective, the optimization is made in terms of profit. But, these two things, the monetary and fiscal policy, end up being uncoordinated.
@@
Another way of understanding what we have done is to recognize that previous economic theory has been built from advanced propositions, not from theorems proved from basic principles, that is not built up from a foundation. These propositions have been either neoclassical capitalism (Keynesianism for fiscal demand-side policy, and Monetarism for monetary supply-side policy) relying on resistive control, or socialism (Marxian Economics, Marxism, Modern Monetary Theory, Complexity Economics, and the Economics of Mutuality) with no method of control. Note that resistivity controls investment, that is allocates capital, by commercial banks loaning money with interest, giving a cost to capital. Resistivity controls the money supply by the central bank loaning money with interest to commercial banks. We have built up a new unified economic theory from a foundation of basic principles, without resorting to resistivity to control the economic collective as neoclassical economics does. Instead, we have used genAI to control the economic collective, yielding a well controlled and high performing socialism. In the case that resistive control is used, a low performing neoclassical capitalism is the result.
@@
This constraint is that all of the returns multiplied by both the network and the savings multipliers ($m_{e}S_{0}/T_{I}\sim 10$) less some of the short-term costs (the ones not covered by short-term returns) be greater than zero. The constraint $\Delta M>0$ is that the increase in the demand for the electronic currency be greater than zero. Just so that it is greater than zero, there will be an increase in the value of the currency and no inflation. The difference between the two constraints is stark – short-term returns less all short-term costs versus all of the returns times $m_{e}S_{0}/T_{I}$ less some of the short-term costs being greater than zero. Much more investment is justified by the $\Delta M>0$ inequality than by the $\text{NPV}>0$ inequality. Looking at the three multipliers in the second line of the expression for NPV in Eq. (157) is very instructive. The term for the DCF, or the upper limit on investment, is the total return $R_{mSR}$ (the upper limit on investment in our theory) times three factors. The first factor is a small factor given by the inverse of the network and savings multipliers $T_{I}/m_{e}S_{0}\sim 0.1$. Remember that these two multipliers come from the fact that an increase of the yearly revenue of a member of the sub-economy is multiplied by both the network effect and the savings effect. The second factor is a term that favors projects with a large profit margin. The third factor is a function that favors projects with a short life time $T_{0}\lesssim 1/\nu$, and quick payback $T_{I}\lesssim 1/\nu$. These are all false incentives, with the first term a false reduction in the potential investment because the network and savings multipliers are not recognized by NPV.
@@
The equity of the business is risk tranched into two securities – the electronic currency (that is, the transactional equity) and the stock (that is, the common equity) of the local economy. The electronic currency is the low risk (that is, small fluctuations), low return, easy to transact (that is, trade) security, and the stock is the high risk (that is large fluctuations), high return, more difficult to transact security. The fluctuations of the electronic currency are further reduced, and the equilibrium stabilized by arbitrage trading of the electronic currency. This arbitrage trading transfers the fluctuations (that is, risk) from the electronic currency to the stock. The fluctuations in the electronic currency are driven by a lack of liquidity in the electronic currency, that is the counter party to the trade not being there at the time of the trade. The arbiter steps in and makes the trade, then makes the opposite trade when the counter party arrives. As the local economy gets larger, the liquidity increases and the required liquid reserve ratio decreases, as one over the square root of the size of the local economy. The arbiter buys for a small amount less than the true currency price and sells for a small amount more than the true currency price. The difficulty is knowing the true currency price. When the arbitrage trading is done according to Eq. (100), the arbiter will always buy low and sell high, creating the financial heat pump described in Sec. VII. The risk tranching is a consequence of the backing (that is, securitization) of the electronic currency by tangible assets whose “replacement” costs can be accounted by transactions (or equivalently the transactional monetary demand given in Eq. (16)), and the backing of the stock by predominately intangible assets whose value can not be accounted for (that is, things like know-how, show-how, value-in-use, goodwill, and brand value). It is essential that when these two securities are traded by the public, that the public have complete real-time on-line visibility of the finances of the local economy. This allows the public to correctly value the securities.
@@
The bottom line is that electronic currency leveraged financing allows more effective operation of sub-economies and more appropriate levels of fiscal investment. This is supercharged by approaching monetary and fiscal policy as a question of financial systems control based on innovative genAI methods of system characterization (i.e., risk estimation) and system control. The economy is operated at the point of maximum sustainable virtuous economic activity, that is at the point of economic and social prosperity. This is in contrast to conventional resistive control. In this case, the supply of capital is controlled by the central bank interest rate (raise the rate to reduce the supply and lower the rate to increase the supply), and the distribution of capital is rationed by the cost of capital, the interest rate at which it is loaned. The economy is operated at the point of maximum short term profit, that is at the point of maximum economic and social exploitation.
@@
The leader has a choice of two ethos: the aesthetic or the unaesthetic. The choice is between a point of global aesthetic maximum where everyone wins or a point of global aesthetic minimum (the unaesthetic) where only a few people win and everyone else looses. This is a choice of the future or the religious eskaton. There are two ways that a strategic player can play a game against a reactionary opponent. The strategic player can lead his opponent to play the game very poorly leading to quick and decisive defeat of the opponent where the strategic player always wins. This is a choice of the unaesthetic, an ugly boring game with the misery of the opponent maximized. The strategic player will feel that the game is boring and unrewarding, too. In contrast, the strategic player can lead his opponent to playing well resulting in a closely matched interesting game where there is a significant chance that the opponent will win. This is a choice of the aesthetic, a beautiful engaging game with the enjoyment of both participants, as well as the spectators, maximized. This is the beautiful game of football of which Brazilian fans speak, the “Joga Bonito”. It is not whether one wins or looses, it is how one plays the game. The best football match ever played was the 2022 World Cup final between Argentina and France. It does not matter that Argentina won the match. France could have easily won with a slightly different trajectory of the ball off of the foot of Mbappé late in the match. What made it the most beautiful match ever was both Messi and Mbappé playing the best football of their career against each other. They both elevated the play of each other to a perfectly matched level of performance never seen before.
@@
In summary, this paper presented a new unified economic theory that relaxed the capitalistic diffusive approximation made for control so that there is coordinated, yet distributed, monetary and fiscal (print/grant investment) policy made to maximize GDP, not DCF. This is a socialist, not a capitalistic metric, where the economic system is treated as a web (multiscale graph) of economic communes. The local microeconomics is treated like the global macroeconomics. The viscous forecasting and control is replaced by a Generative Artificial Intelligence (genAI) forecasting and control. The control is distributed (that is, democratic), not centralized (that is autocratic). Monetarism argues for only an autocratic (centralized) capitalistic (profit based) monetary control, and Keynesianism argues for only an autocratic capitalistic fiscal (borrow/spend investment) control. This new theory argues for coordinated democratic socialist monetary and fiscal control. It does have philosophical connections to African Ubuntu, the Economics of Mutuality, Modern Monetary Theory, Marxian Economics, Complexity Economics, and religious philosophy. It also makes the connection between genAI and the physics of economic collective systems. The result is an economic system that maximizes the good of society, not an economic system that maximally exploits society.